HIGH RESOLUTION INFRARED SPECTROSCOPY OF THE ν_{5} AND ν_{7} BANDS OF HEXAFLUOROETHANE

\underline{THOMAS A. BLAKE}, STEVEN W. SHARPE, Battelle/Pacific Northwest National Laboratory, P. O. Box 999, MS K3- 58, Richland, WA 99352; STEVEN C. STONE, LAURA A. PHILIPS, Department of Chemistry, Cornell University, Ithaca, NY 14853-1301.

The ν_{5} and ν_{7} fundamental bands of hexafluoroethane were observed using two molecular beam techniques: pulsed slit-jet, diode laser spectroscopy, and continuous slit-jet, Fourier transform spectroscopy. The estimated rotational temperature of both beams is on the order of 10 K. Preliminary analysis of the ν_{5} parallel band shows a regular progression of strong K = 0 transitions in the P, Q, and R-branches; other sub-bands are heavily perturbed. The band origin is at 1117.1098(5) cm^{−1}, B′′ = 0.06140(9) cm^{−1}, and B′ = 0.061641 cm^{−1} for the prolate symmetric top molecule. The ν_{7} perpendicular band is centered at 1253 cm^{−1} and shows a congested series of ^{R}Q and ^{P} Q sub-branches. Spectra from both spectrometers and a detailed ro-vibrational analysis will be presented for both bands.